# A Point-Based Algorithm for Distributional Reinforcement Learning in Partially Observable Domains

### v1

@@ Abstract
In many real-world planning tasks, agents must tackle uncertainty about the environment’s state and variability in the outcomes of any chosen policy. We address both forms of uncertainty as a first step toward safer algorithms in partially observable settings. Specifically, we extend Distributional Reinforcement Learning (DistRL)—which models the entire return distribution for fully observable domains—to Partially Observable Markov Decision Processes (POMDPs), allowing an agent to learn the distribution of returns for each conditional plan. Concretely, we introduce new distributional Bellman operators for partial observability and prove their convergence under the supremum p-Wasserstein metric. We also propose a finite representation of these return distributions via $\psi$-vectors, generalizing the classical $\alpha$-vectors in POMDP solvers. Building on this, we develop Distributional Point-Based Value Iteration (DPBVI), which integrates $\psi$-vectors into a standard point-based backup procedure—bridging DistRL and POMDP planning. By tracking return distributions, DPBVI naturally enables risk-sensitive control in domains where rare, high-impact events must be carefully managed. We provide source code to foster further research in robust decision-making under partial observability.
@@ 1 Introduction
Reinforcement Learning (RL) traditionally maximizes expected returns, treating each future reward distribution as a scalar. However, growing evidence shows that modeling the full return distribution, rather than just its mean, can yield robust policies, better exploration, and richer theoretical insights (Bellemare et al., 2023). This distributional perspective, referred to as Distributional Reinforcement Learning (DistRL), captures variability and higher-order statistics of returns, providing agents with more profound information about environment stochasticity.
@@
Beyond maximizing expected returns, many real-world settings require risk-sensitive planning, where limiting the probability of adverse or catastrophic outcomes is crucial. In partially observable domains, the stakes are even higher because the agent’s uncertainty about the state can exacerbate the impact of rare but high-cost events. For instance, a slight misjudgment of the patient’s condition or the vehicle’s surroundings in healthcare or autonomous navigation can lead to undesirable or dangerous consequences. Incorporating risk considerations into POMDP planning explicitly accounts for such worst-case scenarios, making policies more conservative when high losses are possible and balancing exploration against the need to avoid significant harm. Establishing a foundational framework for distributional methods in POMDPs, as we do here, is thus a necessary first step toward risk-sensitive control in partially observable domains.
@@
  • Distributional Point-Based Value Iteration (DPBVI). We adapt Point-Based Value Iteration (PBVI) to the distributional setting, yielding DPBVI, which uses $\psi$-vectors instead of $\alpha$-vectors. Under risk-neutral objectives, DPBVI converges to the same policy as PBVI yet provides a foundation for risk-sensitive extensions by maintaining full return distributions.
Alongside these theoretical results, we release our source code [^1] Omitted for double-blind review. to facilitate further exploration of risk-aware approaches in partially observable domains. Thus, our work unifies partial observability and DistRL, enabling distributionally robust decision-making where state uncertainty and return variability are central considerations.
@@ 2.2 Distributional Reinforcement Learning
Traditionally, the value function models the return expectation. DistRL, instead, aims to learn the compound distribution of the returns $Z^{\pi}(s)$ sourced to randomness in 1. reward $R$ 2. transition $P^{\pi}$, and 3. distribution of the next-state value $Z(S^{\prime})$. Let $\widetilde{\mathcal{T}}$ be the distributional Bellman operator such that
@@
The distributional Bellman operator has been shown to be a $\gamma$-contraction under the supremum $p$-Wasserstein metric space $\forall p\in[1,\infty]$ (Bellemare et al., 2023). The distributional Bellman optimality operator $\widetilde{\mathcal{T}}^{*}$, which selects actions that maximize the expected return, similarly converges to the fixed point under the supremum $p$-Wasserstein metric space $\forall p\in[1,\infty]$ (Bellemare et al., 2023) if there is a unique optimal policy and a mean-preserving distribution representation $\mathcal{F}$ is used (e.g., a categorical representation). In the case of the existence of multiple optimal policies, however, convergence isn’t guaranteed because differing distributions can be similar in expectation.
@@ 3.1 The Partially Observable Distributional Bellman Operators
Proposition 1.
For any POMDP $\mathcal{M}$ with rewards bounded by $[R_{\text{min}},R_{\text{max}}]$ the partially observable distributional Bellman operator is a contraction on $\mathfrak{P}(\mathbb{R})^{\Delta}$ in the supremum p-Wasserstein distance $\bar{w}_{p},\forall p\in[1,\infty),\forall\gamma\in[0,1)$. That is,
@@
$\forall\eta,\eta^{\prime}\in\mathfrak{P}(\mathbb{R})^{\mathcal{S}}$.
Proof.
To prove Proposition 1, we lift $\mathcal{M}$ to an equivalent, fully observable belief MDP $\widetilde{\mathcal{M}}$, where belief transitions are defined by $\tau(b,a,o^{\prime})$ with probability $P(o^{\prime}\mid b,a)=\sum_{s,s^{\prime}\in\mathcal{S}}b(s)\,T(s,a,s^{\prime})\,\Omega(o^{\prime}\mid s^{\prime},a)$. Let $p\in[1,\infty),\gamma\in[0,1)$ be fixed.
$w_{p}^{p}\left((\widetilde{\mathcal{T}}^{\pi}_{PO}Z_{0})(b),(\widetilde{\mathcal{T}}^{\pi}_{PO}Z_{1})(b)\right)$ $\overset{(a)}{\leq}\mathbb{E}_{\pi}\left[|(R+\gamma Z_{0}(B^{\prime}))-(R+\gamma Z_{1}(B^{\prime}))|^{p}\mid B=b\right]$
=$\overset{(b)}{=}\gamma^{p}\mathbb{E}\left[|Z_{0}(B^{\prime})-Z_{1}(B^{\prime})|^{p}\mid B=b\right]$
$\overset{(c)}{\leq}\gamma^{p}\int_{b^{\prime}\in\Delta}P^{\pi}(B^{\prime}=b^{\prime}\mid B=b)\mathbb{E}\left[|Z_{0}(b^{\prime})-Z_{1}(b^{\prime})|^{p}\right]$
$\overset{(d)}{\leq}\sup_{b^{\prime}\in\Delta}\mathbb{E}\left[|Z_{0}(b^{\prime})-Z_{1}(b^{\prime})|^{p}\right]$
=$\overset{(e)}{=}\gamma^{p}\bar{w}^{p}_{p}(Z_{0},Z_{1})$
(a) follows from the Wasserstein distance being defined as a minimum over couplings, (b) results from algebraically adjusting the expectation, (c) follows from the fact that the belief transitions and the return random variables are treated independently in the coupling, and the belief space being continuous necessitates integration, (d) results from the maximum of nonnegative quantities being at least as great as their weighted average, and (e) follows since $(Z_{0},Z_{1})$ is defined as an optimal coupling. ∎
Proposition 1 follows from the standard $\gamma$-contraction properties of $\widetilde{\mathcal{T}}^{\pi}$. Because $\widetilde{\mathcal{T}}^{\pi}_{PO}$ is a contraction mapping, by the Banach Fixed-Point Theorem, $\exists!\eta^{*}\in\mathfrak{P}$ s.t. $\eta^{*}$ is a fixed point.
Proposition 2.
Let $\widetilde{\mathcal{T}}^{*}_{PO}$ be the partially observable distributional Bellman optimality operator for a POMDP $\mathcal{M}$ with bounded rewards and let $p\in[1,\infty]$. Suppose there is a unique optimal policy $\pi^{*}$ and, for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the reward distribution is $P_{\mathfrak{R}}(\cdot\mid s,a)$. For any initial return-distribution function $\eta_{0}\in\mathfrak{P}(\mathbb{R})$, the sequence of iterates
@@
Proof.
We again view the POMDP $\mathcal{M}$ through its equivalent fully observable belief MDP $\widetilde{\mathcal{M}}$, where each belief $b\in\Delta$ in $\mathcal{M}$ becomes a state of $\widetilde{\mathcal{M}}$. Noting this setup is directly analogous to the greedy selection rule $\mathcal{G}$ in Bellemare et al. (2023, Theorem 7.9), we can directly apply this theorem and conclude that the iterates converge to the unique, optimal return distribution $\eta^{\pi^{*}}$. ∎
@@
Consider a mean-preserving projection $\Pi_{\mathcal{F}}$ for some representation $\mathcal{F}$. Suppose $\Pi_{\mathcal{F}}$ is a nonexpansion in the supremum $p$-Wasserstein distance, and let the finite domain $\mathfrak{P}_{\mathcal{R}}(\mathbb{R})$ be closed under $\Pi_{\mathcal{F}}$. If there is a unique optimal policy $\pi^{*}$ then, under the conditions of Proposition 2, the sequence
@@
Proposition 2 states if there is a unique optimal policy, repeatedly applying $\widetilde{\mathcal{T}}^{*}_{PO}$ will converge to the optimal return distribution. Corollary 1 establishes that if we approximate the return distribution after each greedy update, the resulting approximate iteration still converges to the same fixed point, provided there is a unique optimal policy.
@@ 3.2 $\psi$-vectors
Proposition 3.
Under the assumptions of Proposition 2 and Corollary 1, the optimal distributional value function $Z^{*}(b)$ can be represented as a finite set of $\psi$-vectors $\Psi$, $\forall a\in\mathcal{A}$, such that for any belief state $b\in\Delta$,
@@
Proof.
=$\overset{(a)}{=}\max_{\Psi\in\Gamma^{*}}\langle\Psi,b\rangle$
=$\overset{(b)}{=}\max_{\alpha\in\Gamma^{*}}\alpha\cdot b$
=$\overset{(c)}{=}V^{*}(b)$
By Proposition 3, $Z^{*}$ inherits the PWLC structure of classical POMDPs, but within the Wasserstein metric space. Crucially, this representation remains finite in the risk-neutral case, ensuring computational tractability—just as a finite set of $\alpha$-vectors suffices to represent the classical optimal value function $V^{*}$.
@@ 4 Distributional Point-Based Value Iteration
By Proposition 3, a finite set of these $\psi$-vectors suffices to represent the optimal distributional value function $Z^{*}$ in the risk-neutral setting, just as $\alpha$-vectors suffice in classical POMDPs. We represent return distributions using the categorical distribution parameterization, which is mean-preserving when its support spans all possible returns (Bellemare et al., 2023). This ensures that the expected returns align with the classical POMDP value function, preserving the key convergence guarantees from Corollary 1 while operating in the space of return distributions.
@@ 4.1 The DBPVI Backup
Because we manage distributions instead of scalars, each backup is slower by a factor of $|M|$, where $|M|$ is the number of bins in each categorical distribution. Concretely, one DPBVI backup runs in $\mathcal{O}(|\mathcal{S}|\;|\mathcal{A}|\;|\Gamma^{t-1}|\;|\mathcal{O}|\;|\mathcal{B}|\;|M|)$, compared to $\mathcal{O}(|\mathcal{S}|\;|\mathcal{A}|\;|V_{t-1}|\;|\mathcal{O}|\;|\mathcal{B}|)$ for PBVI. Nonetheless, DPBVI preserves the core PBVI structure—belief set, backup iteration, and finite policy representation—while enabling distributional modeling of returns under patial observability.

### v2

# Reinforcement Learning under State and Outcome Uncertainty: A Foundational Distributional Perspective

@@ Abstract
In many real-world planning tasks, agents must tackle uncertainty about the environment’s state and variability in the outcomes of any chosen policy. We address both forms of uncertainty as a first step toward safer algorithms in partially observable settings. Specifically, we extend Distributional Reinforcement Learning (DistRL)—which models the entire return distribution for fully observable domains—to Partially Observable Markov Decision Processes (POMDPs), allowing an agent to learn the distribution of returns for each conditional plan. Concretely, we introduce new distributional Bellman operators for partial observability and prove their convergence under the supremum $p$-Wasserstein metric. We also propose a finite representation of these return distributions via $\psi$-vectors, generalizing the classical $\alpha$-vectors in POMDP solvers. Building on this, we develop Distributional Point-Based Value Iteration (DPBVI), which integrates $\psi$-vectors into a standard point-based backup procedure—bridging DistRL and POMDP planning. By tracking return distributions, DPBVI lays the foundation for future risk-sensitive control in domains where rare, high-impact events must be carefully managed. We provide source code to foster further research in robust decision-making under partial observability.
@@ 1 Introduction
Reinforcement Learning (RL) traditionally maximizes expected returns, treating each future reward distribution as a scalar. However, growing evidence suggests that modeling the full return distribution, rather than just its mean, can yield robust policies, better exploration, and richer theoretical insights (Bellemare et al., 2023). This distributional perspective, referred to as Distributional Reinforcement Learning (DistRL), captures variability and higher-order statistics of returns, providing agents with more profound information about environment stochasticity.
@@
Beyond maximizing expected returns, many real-world settings require risk-sensitive planning, where limiting the probability of adverse or catastrophic outcomes is crucial. In partially observable domains, the stakes are even higher because the agent’s uncertainty about the state can exacerbate the impact of rare but high-cost events. For instance, a slight misjudgment of the patient’s condition or the vehicle’s surroundings in healthcare or autonomous navigation respectively can lead to undesirable or dangerous consequences. Incorporating risk considerations into POMDP planning explicitly accounts for such worst-case scenarios, making policies more conservative when high losses are possible and balancing exploration against the need to avoid significant harm. Establishing a foundational framework for distributional methods in POMDPs, as we do here, is thus a necessary first step toward risk-sensitive control in partially observable domains. Although we do not formally address risk-sensitive objectives, modeling full return distributions under partial observability is a necessary prerequisite for future work in this direction.
@@
  • Distributional Point-Based Value Iteration (DPBVI). We adapt Point-Based Value Iteration (PBVI) to the distributional setting, yielding DPBVI, which uses $\psi$-vectors instead of $\alpha$-vectors. Under risk-neutral objectives, DPBVI converges to the same policy as PBVI yet lays the groundwork for risk-sensitive extensions by maintaining full return distributions.
Alongside these theoretical results, we release our source code [^1] Source code available at github.com/lpreuettUW/distributional_point_based_value_iteration. to facilitate further exploration of risk-aware approaches in partially observable domains. Thus, our work unifies partial observability and DistRL, enabling distributionally robust decision-making where state uncertainty and return variability are central considerations.
@@ 2.2 Distributional Reinforcement Learning
Traditionally, the value function models the return expectation. DistRL, instead, aims to learn the compound distribution of the returns $Z^{\pi}(s)$ sourced to randomness in (1) reward $R$ (2) transition $P^{\pi}$, and (3) distribution of the next-state value $Z(S^{\prime})$. Let $\widetilde{\mathcal{T}}$ be the distributional Bellman operator such that
@@
The distributional Bellman operator has been shown to be a $\gamma$-contraction under the supremum $p$-Wasserstein metric space $\forall p\in[1,\infty]$ (Bellemare et al., 2023). The distributional Bellman optimality operator $\widetilde{\mathcal{T}}^{*}$, which selects actions that maximize the expected return, similarly converges to the fixed point under the supremum $p$-Wasserstein metric space $\forall p\in[1,\infty]$ (Bellemare et al., 2023) if there is a unique optimal policy and a mean-preserving distribution representation $\mathcal{F}$ is used (e.g., a categorical distribution representation). In the case of the existence of multiple optimal policies, however, convergence isn’t guaranteed because differing distributions can be similar in expectation.
@@ 3.1 The Partially Observable Distributional Bellman Operators
We denote by $\mathfrak{P}_{p}(\mathbb{R})$ the set of Borel probability measures on $\mathbb{R}$ with finite $p$-th moment, ensuring that the $p$-Wasserstein distance is well-defined. The space $\mathfrak{P}_{p}(\mathbb{R})^{\Delta}$ then denotes the set of functions $\eta\ :\Delta\rightarrow\mathfrak{P}_{p}(\mathbb{R})$, where $\Delta$ is the belief space of the POMDP.
Theorem 1.
For any POMDP $\mathcal{M}$ with rewards bounded by $[R_{\text{min}},R_{\text{max}}]$ the partially observable distributional Bellman operator is a contraction on $\mathfrak{P}_{p}(\mathbb{R})^{\Delta}$ in the supremum $p$-Wasserstein distance $\bar{w}_{p},\forall p\in[1,\infty),\forall\gamma\in[0,1)$. That is,
@@
$\forall\eta,\eta^{\prime}\in\mathfrak{P}_{p}(\mathbb{R})^{\Delta}$.
See Appendix A.1 for proof.
Theorem 1 follows from the standard $\gamma$-contraction properties of $\widetilde{\mathcal{T}}^{\pi}$. Because $\widetilde{\mathcal{T}}^{\pi}_{PO}$ is a contraction mapping, by the Banach Fixed-Point Theorem, there exists a unique $\eta^{*}\in\mathfrak{P}_{p}(\mathbb{R})^{\Delta}$ such that $\eta^{*}$ is a fixed point.
Theorem 2.
Let $\widetilde{\mathcal{T}}^{*}_{PO}$ be the partially observable distributional Bellman optimality operator for a POMDP $\mathcal{M}$ with bounded rewards and let $p\in[1,\infty]$. Suppose there is a unique optimal policy $\pi^{*}$ and, for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the reward distribution is $P_{\mathfrak{R}}(\cdot\mid s,a)$. For any initial return-distribution function $\eta_{0}\in\mathfrak{P}_{p}(\mathbb{R})$, the sequence of iterates
@@
Consider a mean-preserving projection $\Pi_{\mathcal{F}}$ for some representation $\mathcal{F}$. Suppose $\Pi_{\mathcal{F}}$ is a nonexpansion in the supremum $p$-Wasserstein distance, and let the finite domain $\mathfrak{P}_{p}(\mathbb{R})^{\Delta}$ be closed under $\Pi_{\mathcal{F}}$. If there is a unique optimal policy $\pi^{*}$ then, under the conditions of Theorem 2, the sequence
@@
These results build on the convergence guarantees of the distributional Bellman optimality operator in Bellemare et al. (2023, Theorem 7.9), which assumes a unique optimal policy and analyzes the tabular MDP setting. Our proof adapts these arguments to the belief MDP by ensuring the required continuity and compactness conditions hold, allowing the result to generalize to partially observable settings (see Appendix A.2). Bellemare et al. note that the unique optimal policy assumption is necessary for contraction guarantees but they empirically observe that algorithms like C51 often perform well even when this condition may not strictly hold Bellemare et al. (2023).
Theorem 2 states if there is a unique optimal policy, repeatedly applying $\widetilde{\mathcal{T}}^{*}_{PO}$ will converge to the optimal return distribution. Corollary 1 establishes that if we approximate the return distribution after each greedy update, the resulting approximate iteration still converges to the same fixed point, provided there is a unique optimal policy.
@@ 3.2 $\psi$-vectors
Theorem 3.
Under the assumptions of Theorem 2 and Corollary 1, the optimal distributional value function $Z^{*}(b)$ can be represented as a finite set of $\psi$-vectors $\Psi$, $\forall a\in\mathcal{A}$, such that for any belief state $b\in\Delta$,
@@
See Appendix A.3 for proof.
By Theorem 3, $Z^{*}$ inherits the PWLC structure of classical POMDPs, but within the Wasserstein metric space. Crucially, this representation remains finite in the risk-neutral case, ensuring computational tractability—just as a finite set of $\alpha$-vectors suffices to represent the classical optimal value function $V^{*}$.
@@ 4 Distributional Point-Based Value Iteration
By Theorem 3, a finite set of these $\psi$-vectors suffices to represent the optimal distributional value function $Z^{*}$ in the risk-neutral setting, just as $\alpha$-vectors suffice in classical POMDPs. We represent return distributions using the categorical distribution parameterization, which is mean-preserving when its support spans all possible returns (Bellemare et al., 2023). This ensures that the expected returns align with the classical POMDP value function, preserving the key convergence guarantees from Corollary 1 while operating in the space of return distributions.
@@ 4.1 The DBPVI Backup
Because we manage distributions instead of scalars, each backup is slower by a factor of $|M|$, where $|M|$ is the number of bins in each categorical distribution. Concretely, one DPBVI backup runs in $\mathcal{O}(|\mathcal{S}|\;|\mathcal{A}|\;|\Gamma^{t-1}|\;|\mathcal{O}|\;|\mathcal{B}|\;|M|)$, compared to $\mathcal{O}(|\mathcal{S}|\;|\mathcal{A}|\;|V_{t-1}|\;|\mathcal{O}|\;|\mathcal{B}|)$ for PBVI. Nonetheless, DPBVI preserves the core PBVI structure—belief set, backup iteration, and finite policy representation—while enabling distributional modeling of returns under partial observability.
@@ 5 Experimental Results
We empirically validate that DPBVI recovers the same value function as PBVI under risk-neutral objectives, as predicted by theory. Our focus is not on performance or efficiency, but on whether distributional backup operations preserve correctness relative to standard scalar PBVI.
We evaluate DPBVI and PBVI in two small POMDP domains: (1) MiniGrid DoorKey (5×5), where the agent must reach a goal behind a locked door using a key, and (2) a Two-State Noisy-Sensor domain with binary state transitions and stochastic observations. See Appendix B for full environment and implementation details.
Table 1 reports runtime and iteration counts for both methods. As expected, DPBVI incurs greater computational cost due to distributional operations but converges in the same number of iterations as PBVI. To assess value function agreement, we compare the expected values of DPBVI’s $\psi$-vectors with PBVI’s scalar values across belief points. Figure 1 shows the maximum relative error per iteration. Errors remain very small, confirming that DPBVI tracks PBVI’s value function closely.
We further test for compounding error by tightening the convergence threshold in the Two-State Noisy-Sensor environment (Figure 2). The relative error initially increases, then stabilizes and decreases, suggesting discrepancies stem from categorical projection, floating-point limits, and bootstrapping rather than theoretical flaws. These findings confirm that DPBVI is a correct extension of PBVI under risk-neutral assumptions and establish a foundation for future risk-sensitive variants.
6 Conclusion
In this paper, we presented the first comprehensive study of Distributional Reinforcement Learning in partially observable domains. Our main theoretical result showed that the partially observable distributional Bellman operator $\widetilde{\mathcal{T}}_{PO}^{\pi}$ and its optimality counterpart $\widetilde{\mathcal{T}}^{*}_{PO}$ are $\gamma$-contractions in the supremum $p$-Wasserstein metric ($1\leq p<\infty$). These results extend the well-known contraction properties of distributional Bellman operators in fully observable settings to POMDPs.
As part of our analysis, we introduced $\psi$-vectors as the distributional analogs of $\alpha$-vectors, proving that under risk-neutral control, the optimal distributional value function in a POMDP can be represented by a finite collection of these $\psi$-vectors while preserving PWLC in the Wasserstein metric space. This representation underlines how a distributional perspective naturally generalizes classical POMDP theory.
Building on these foundations, we described DPBVI, a risk-neutral algorithm that applies the distributional Bellman optimality operator to a finite set of belief points. This extension parallels the classical PBVI procedure yet replaces scalar $\alpha$-vectors with distributional $\psi$-vectors, thereby capturing richer information about uncertainty in returns. Experiments (see Section 5 and Appendix B) show that DPBVI recovers the same solution as PBVI under risk-neutral control, confirming that the distributional approach does not compromise solution quality in the risk-neutral regime.
While our empirical and theoretical results focus solely on the risk-neutral setting, this work enables future exploration of risk-sensitive approaches by extending the distributional perspective to partially observable domains. Capturing full return distributions under uncertainty lays the groundwork for future work on risk-sensitive control. We hope these developments foster further exploration of DistRL in practical domains where both state uncertainty and variability in outcomes are central concerns.
That said, while our experiments validate theoretical correctness, they are limited to small synthetic environments and do not assess scalability or empirical performance in large-scale settings. Future work may explore practical extensions of DPBVI to high-dimensional problems and risk-sensitive criteria.
Acknowledgments
The authors thank Ankur Teredesai and Kevin Jamieson for their helpful feedback on early drafts of this work. Muhammad Aurangzeb Ahmad’s work was supported by the National Institutes of Health, USA (NIH grant T32 GM 121290).
Appendix A Proofs
A.1 Proof of Theorem 1
See 1
Proof.
We prove contraction of the partially observable distributional Bellman operator $\widetilde{\mathcal{T}}^{\pi}_{PO}$ by lifting the POMDP $\mathcal{M}$ to its equivalent fully observable belief MDP $\widetilde{\mathcal{M}}$, where states are beliefs $b\in\Delta$, actions are $a\in\mathcal{A}$, and transitions are governed by the belief update $\tau(b,a,o^{\prime})$ with probability
$P(o^{\prime}\mid b,a)=\sum_{s,s^{\prime}}b(s)T(s,a,s^{\prime})\Omega(o^{\prime}\mid s^{\prime},a).$
Assume the following:
  • $\Delta$ is a Borel space (as it is a subset of a finite-dimensional simplex).
  • $\mathcal{A}$ is finite.
  • The reward function $R(b,a):=\mathbb{E}_{s\sim b}[R(s,a)]$ is bounded.
  • The belief transition kernel $\tau(\cdot\mid b,a)$ is measurable.
  • The policy $\pi:\Delta\to\mathcal{A}$ is fixed and measurable.
Let $\eta,\eta^{\prime}\in\mathfrak{P}_{p}(\mathbb{R})^{\Delta}$ be two distribution-valued value functions. Define the supremum $p$-Wasserstein metric as:
$\bar{w}_{p}(\eta,\eta^{\prime}):=\sup_{b\in\Delta}W_{p}(\eta(b),\eta^{\prime}(b)),\quad\forall p\in[1,\infty).$
Fix $b\in\Delta$ and let $(G_{b},G_{b}^{\prime})$ be an optimal coupling of $\eta(b)$ and $\eta^{\prime}(b)$. Define:
$\tilde{G}_{b}:=R(b,\pi(b))+\gamma G_{b^{\prime}},\quad\tilde{G}_{b}^{\prime}:=R(b,\pi(b))+\gamma G^{\prime}_{b^{\prime}},$
where $b^{\prime}\sim\tau(\cdot\mid b,\pi(b))$ is the next belief under the belief transition kernel.
Then:
$W_{p}(\widetilde{\mathcal{T}}^{\pi}_{PO}\eta(b),\widetilde{\mathcal{T}}^{\pi}_{PO}\eta^{\prime}(b))\leq\mathbb{E}_{b^{\prime}}\left[W_{p}(\eta(b^{\prime}),\eta^{\prime}(b^{\prime}))\right]\leq\gamma\bar{w}_{p}(\eta,\eta^{\prime}).$
Taking the supremum over $b\in\Delta$, we conclude:
$\bar{w}_{p}\left(\widetilde{\mathcal{T}}^{\pi}_{PO}\eta,\widetilde{\mathcal{T}}^{\pi}_{PO}\eta^{\prime}\right)\leq\gamma\bar{w}_{p}(\eta,\eta^{\prime}),$
as desired. ∎
A.2 Proof of Theorem 2
See 2
Proof.
We lift the POMDP $\mathcal{M}$ to its equivalent belief MDP $\widetilde{\mathcal{M}}$, where each belief $b\in\Delta$ serves as a fully observable state. The transition dynamics $\tau(b,a,o^{\prime})$ of the belief MDP are Markovian, and the reward function is defined as $\mathcal{R}(b,a):=\mathbb{E}_{s\sim b}[\mathcal{R}(s,a)]$, which is bounded by assumption.
From this point forward, we use $b\in\Delta$ to denote the state of the belief MDP. We assume the following conditions hold in the belief MDP:
  • The belief state space $\Delta$ is compact and convex
  • The reward function $\mathcal{R}(b,a)$ is continuous in $b\;\forall a\in\mathcal{A}$
  • The belief update function $\tau(b,a,o^{\prime})$ is continuous in $b$ for all $a\in\mathcal{A}$ and $o^{\prime}\in\mathcal{O}$
  • The action space $\mathcal{A}$ is finite
  • The policy space $\Pi$ is fixed and measurable
  • There exists a unique optimal policy $\pi^{*}$
Let $\eta^{\pi^{*}}$ denote the return distribution induced by the optimal policy. The partially observable distributional Bellman optimality operator $\widetilde{\mathcal{T}}^{*}_{PO}$ can be interpreted as a greedy update rule over return distributions in the belief MDP.
Define $Q_{\eta}(b,a):=\mathbb{E}_{Z\sim\eta(b,a)}[Z]$ as the expected return under the return distribution $\eta$. The action gap at state $b\in\Delta$ is defined as
$\text{GAP}(Q_{\eta},b):=\min\{Q_{\eta}(b,a^{*})-Q_{\eta}(b,a):a^{*},a\in\mathcal{A},a^{*}\neq a,Q_{\eta}(b,a^{*}):=\max_{a^{\prime}\in\mathcal{A}}Q_{\eta}(b,a^{\prime})\}$
The global action gap is then defined as
$\text{GAP}:=\inf_{b\in\Delta}\text{GAP}(Q_{\eta},b)$
We now justify that $Q_{\eta}(b,a)$ is continuous in $b$ for all $a\in\mathcal{A}$. The return distribution $\eta(b,a)$ is defined recursively from the reward distribution and belief transitions. Under our assumptions that the reward function $\mathcal{R}(b,a)$ and belief update $\tau(b,a,o^{\prime})$ are continuous in b, it follows that $Q_{\eta}(b,a)=\mathbb{E}_{Z\sim\eta(b,a)}[Z]$ is continuous in $b$ as well. Since the action space is finite, the maximum and second maximum of $\{Q_{\eta}(b,a)\}_{a\in\mathcal{A}}$ are continuous in $b$, so $\text{GAP}(Q_{\eta},b)$ is also continuous in $b$.
By the Extreme Value Theorem, since $\text{GAP}(Q_{\eta},b)$ is continuous on the compact state space $\Delta$, the infimum is attained and equal to the minimum. Since the optimal policy is unique, $\text{GAP}(Q_{\eta},b)>0\;\forall b\in\Delta$, thus $\text{GAP}>0$.
Fix $\epsilon=\dfrac{1}{2}\text{GAP}(Q^{*})$. The standard Bellman optimality operator is known to be a $\gamma$-contraction under the $L^{\infty}$ norm. Consequently, $\exists k\in\mathbb{N}$ such that
$||Q_{\eta_{k}}-Q^{*}||_{\infty}<\epsilon\qquad\forall k\geq K$
For any fixed $b$, let $a^{*}$ be the optimal action in that state. Then for any $a\neq a^{*}$, we have that
$Q_{\eta_{k}}(b,a^{*})$ $\geq Q^{*}(b,a^{*})-\epsilon$
$\geq Q^{*}(b,a)+\text{GAP}(Q^{*})-\epsilon$
$>Q_{\eta_{k}}(b,a)+\text{GAP}(Q^{*})-2\epsilon$
$=Q_{\eta_{k}}(b,a)$
Thus, any greedy selection rule applied to $\eta_{k}$ will yield the optimal policy $\pi^{*}\;\forall k\geq K$. From this point onward, the Bellman updates correspond to evaluation under a fixed policy $\pi^{*}$, and we may invoke Theorem 1 with the initial return distributions $\eta_{0}=\eta_{k}$ to conclude that $\eta_{k}\rightarrow\eta^{\pi^{*}}$. Hence, $\widetilde{\mathcal{T}}^{*}_{PO}$ is a $\gamma$-contraction and converges to the unique fixed point under the supremum $p$-Wasserstein metric.
A.3 Proof of Theorem 3
See 3
Proof.
=$\overset{(a)}{=}\max_{\Psi\in\Gamma^{*}}\langle\Psi,b\rangle$
=$\overset{(b)}{=}\max_{\alpha\in\Gamma^{*}}\alpha\cdot b$
=$\overset{(c)}{=}V^{*}(b)$
Appendix B Detailed Experimental Results